If a=10, b =13 and the angle between a and b is 59.5 degree, what is the scalar product of a and b
1
Expert's answer
2019-09-16T10:49:47-0400

Given that "a=10,b=13, \theta=59.5\degree."

Scalar product


"\bm{a}\cdot\bm{b}=ab\cos{\theta}""\bm{a}\cdot\bm{b}=10 \cdot13\cdot\cos{59.5\degree}\approx65.98"